OpenStudy (anonymous):

simplify by removing factors of 1 14x^2-686 / 4x^2-196

jhonyy9 (jhonyy9):

4x2 -196 =(2x-14)(2x+14)

jhonyy9 (jhonyy9):

2x+14 * 7x-49 = 14x2 -98x +98x -686 =14x2 -686

jhonyy9 (jhonyy9):

(2x+14)(7x+49)/(2x+14)(2x-14)=(7x+49)/(2x-14)=7(x+7)/(2(x-7))